Handwrite in text plz not in pic since its hard to read from and plz dont copy answers that were answered before least 2-3 paragraphs

1. Can pollution ever make us better off? How do we know? Should we aim to eliminate all pollution? If not, what should our goal be? Defend your answer

2. When thinking about types of goods, what does rivalry mean? What does excludability mean? What are the four categories of goods we can identify using those attributes, and what is an example of a good in each category?

3. Flavor Flav has released a new exercise DVD for his loyal fans. If he charges $20 he expects to sell 10,000 copies; if he charges $30, he expects to sell 8,000 copies. Calculate the elasticity of demand for this product between those two points. Is this demand elastic or inelastic?
Answer #1

Answer 1.)

Pollution is a type of negative externality which affects the health of the individual and this leads to the overall decline in the Welfare of the society.

Economically, the pollution leads to inefficiency as it increases the cost in terms of health and depletion of resources etc. which makes the growth of a society unsustainable in the future
Pollution leads to various respiratory diseases in human beings and also the other creatures in danger and imbalances the natural biodiversity of Earth.
Though we cannot eliminate the whole pollution but at least we can strive to reduce the level of the pollution in an economy and bring it to the level which is tolerable.

We could further impose some emission regulations on the firms either in the form of tax or unit capping so that the pollution level remains at the lowest level.


Answer 2.)
Non rivalry is the property of a public good which means that it does not cost any extra money when the consumption of public good increases by a unit.
On the other hand, the non- excludability of a public means that it is not possible to exclude any individual from using the public good.

4 categories the can be identified using the property of rivalry and excludability are as follows:

(i) Non-rival and excludable:
This property refers to the goods which are non-rival and excludable in nature.
Example, private parks to come excludable when entry fees is imposed on the individual.

(ii) Rival and Excludable:
It refers to the private goods which are rival and excludable in nature as an individual is supposed to pay some amount to use these goods.

(iii) Rival and Non-excludable:
For example, the fisheries in the ponds is an example of this type of goods as it is non excludable but when more and more people start using it then it becomes rival.
The fishes in the Ponds start declining due to overuse.

(iv) Non rival and non excludable:
This property relates to the pure public goods which are vavailable to the public without any cost.

Ans 3.)

Initial price = $20

Initial consumption = 10,000 copies

Final price = $30

Final consumption = 8000 copies

Elasticity = - (% change in quantity)/(% change in price)

  

% change in price = (10/20)×100 = 50%

% change in quantity = -(2000/10000)×100 = -20%

Elasticity = -(20/50) = - 0.4

Demand is inelastic as the value "0.4" is less than 1.